# From Voice to Safety: Language AI Powered Pilot-ATC Communication Understanding for Airport Surface Movement Collision Risk Assessment

### v1

@@ 5 Conclusions
In this paper, we introduce a novel Language AI-powered framework for understanding pilot-ATC communication to enhance surface collision risk assessment in ground movement. Our work bridges the gap between traditional surface safety systems (i.e., ASSC) and advanced natural language processing techniques by integrating NER with a surface collision risk model. The proposed approach demonstrates how language-extracted insights from ATC communications can significantly improve surface movement collision risk estimation, as well as providing reference to compliance monitoring.
@@
Our hybrid learning framework lies in the ATC Rule-Enhanced Named Entity Recognition (NER) model, which utilizes domain-specific rules from two ATC manuals regulated by FAA. Experimental results show that integrating these rules substantially improves NER performance across multiple token-level embedding models, with multilingual RoBERTa and BART models achieving the highest F1 scores but scarifying the inference time. The study on time-space complexity trade-off of such hybrid model not only enhances the recognition of critical entities such as flight callsigns, aircraft states, and destination intents but also maintains computational efficiency suitable for near real-time applications.
@@
Our surface collision risk model leverages node-link graph structures of airport layouts and models aircraft taxi speeds with log-normal distributions. Through probabilistic convolution techniques, the model estimates the likelihood of aircraft reaching potential collision nodes simultaneously. Two case studies—one involving a runway collision at Haneda Airport and another a taxiway collision at Hartsfield-Jackson Atlanta International Airport—validated the effectiveness of our methodology. The risk maps generated in both cases accurately highlighted high-risk nodes and demonstrated the practical utility of the proposed framework in real-world scenarios.
5.1 Limitations and Future Work
Our work suffers from several limitations, and several future works are suggested to remedy these concerns,
1.
Further investigation into more accurate NER models is highly recommended. The use of fine-tuned embedding models on aviation-domain corpora, such as Aviation-BERT [61], with a custom classification head (e.g., Bi-LSTM-CRF [60]), has shown improved NER capabilities and can be integrated into our proposed framework seamlessly.
2.
Further investigation and expansion of the ATC rule dictionary is vital for real-world applications, and guarantee the completeness of the proposed method. Significant engineering effort is expected.
3.
Our risk model formulation currently considers collisions between two aircraft only. Extending the model to monitor multi-aircraft surface collision risks is critical for more comprehensive safety monitoring.
4.
The risk model assumes static distributions for taxi speeds. Investigating dynamic, context-aware models that incorporate real-time traffic and weather conditions could improve prediction accuracy.
5.
The taxi speed distributions is assumed to be log-normally distributed in this study but can also be normally distributed for some links. We would suggest a thorough study across more nodes from the node-link graph of a complete taxi time study, potentially grouped based on weight class as well.
6.
Computer vision, as another demonstrated achievement of AI for real-world applications, can contribute to surface movement monitoring by providing timely warnings and cross-verifying aircraft locations with ATC voice communication clearances. By leveraging object detection techniques, both compliance checks and the risk assessment functionalities illustrated in Figure 2 can be enhanced on top of language AI predictions.
7.
The security aspect of the embedding vectors in language AI models are gaining attentions [125, 126]. Differential Privacy (DP) is viewed as a valid method to defend against attacks (i.e., embedding inversions) toward NLP embeddings. A DP fine-tuned embedding model will help increase the safety level of our proposed framework [127].
@@ 5.1 Limitations and Future Work
This work presents a significant step toward harnessing the capabilities of Language AI in aviation safety. By coupling advanced NLP techniques with established risk modeling methodologies, the proposed framework enhances situational awareness and provides a robust tool for mitigating surface collision risks in increasingly complex airport environments. The integration of language understanding into surface movement risk assessment introduces new avenues for enhancing airport safety. By enabling automated detection of miscommunications and deviations from ATC instructions, our approach contributes to proactive incident prevention to enhance aviation safety.

### v2

@@ 5 Conclusions
In this paper, we introduce a novel Language AI-powered framework for understanding pilot-ATC communication to enhance surface collision risk assessment in ground movement. Our work bridges the gap between traditional surface safety systems (i.e., ASSC) and advanced natural language processing techniques by integrating NER with a surface collision risk model. The proposed approach demonstrates how language-extracted insights from ATC communications can significantly improve surface movement collision risk estimation, as well as provide a reference for compliance monitoring.
@@
Our hybrid learning framework lies in the ATC Rule-Enhanced Named Entity Recognition (NER) model, which utilizes domain-specific rules from two ATC manuals regulated by the FAA. Experimental results show that integrating these rules substantially improves NER performance across multiple token-level embedding models, with multilingual RoBERTa and BART models achieving the highest F1 scores but sacrificing the inference time. The study on time-space complexity trade-off of such a hybrid model not only enhances the recognition of critical entities such as flight callsigns, aircraft states, and destination intents but also maintains computational efficiency suitable for near real-time applications.
@@
Our surface collision risk model leverages node-link graph structures of airport layouts and models aircraft taxi speeds with log-normal distributions. Through probabilistic convolution techniques, the model estimates the likelihood of aircraft reaching potential collision nodes simultaneously. Moreover, we propose the real-time risk assessment framework to obtain the collision risk time series at every overlapping node, and demonstrate effectiveness through three real-world case studies. Three case studies demonstrate the effectiveness of our methodology. The risk maps generated in all cases accurately highlighted high-risk nodes and demonstrated the practical utility of the in-time collision risk warning in real-world scenarios.
We acknowledge several limitations and future directions for this research work. First, the NER component can be strengthened by adopting aviation-domain embeddings (e.g., Aviation-BERT [62]) with a tailored classifier such as a Bi-LSTM-CRF head [61]. Second, a broader and better-engineered ATC rule dictionary is needed to ensure completeness and support real-world deployment. Third, our current risk model monitors pairwise surface conflicts only; extending it to multi-aircraft interactions is essential for comprehensive safety assessment. Fourth, the model presently assumes static taxi-speed distributions; developing dynamic, context-aware speed models that reflect real-time traffic and weather should improve predictive accuracy. Fifth, computer vision can augment language-based predictions by detecting and tracking aircraft to cross-check compliance with clearances and enhance the risk-assessment pipeline in Figure 2. Finally, since embedding security is an increasing concern [133, 134], differentially private fine-tuning [135] offers a promising path to mitigate inversion attacks and bolster the safety of the overall framework.
@@
This work presents a significant step toward harnessing the capabilities of Language AI in aviation safety. By coupling advanced NLP techniques with established risk modeling methodologies, this work enhances situational awareness and provides a robust tool for mitigating surface collision risks in increasingly complex airport environments. The integration of language understanding into surface movement risk assessment introduces new avenues for enhancing airport safety. By enabling automated detection of miscommunications and deviations from ATC instructions, our approach contributes to proactive real-time incident prevention to enhance aviation safety.
@@ Acknowledgment
Appendix A Spatiotemporal Risk Error Estimate
This section is a discussion to analysis the approximation error in the spatiotemporal risk formulation. In Equation 14, the linearized equation,
We do a change of variables to normalize by , and set the integral limits to , the non-dimensional ratio of collision radius to distance traveled.
The distribution is computed with a convolution, as it is the sum of the time to reach the previous node from the starting location , and the time to reach the collision location from the previous node . The supremum of the second spatial derivative of may be bounded by using the supremum of the second spatial derivative of .
We perform a change of variables from the time distribution to the log-normal velocity distribution parameterized by and , and take the second derivative with respect to space.
It is apparent that this function may be reorganized as the product of a second-order polynomial with the exponential of another second-order polynomial in terms of . The term in the denominator is kept aside to provide a non-dimensional ratio .
We complete the square and perform a change of variables to leave only a quadratic term in the exponential.
We use known suprema for the individual terms of the expression to bound the supremum for positive .
The dependence of risk error coefficient on taxi speed distribution parameters is shown in Figure 19. Short distances with relatively low speed uncertainty lead to higher error bounds. The upper bound is evaluated at and plugged back into (48). The denominator term is canceled because the second derivative is with respect to the position normalized by .
In conclusion, the approximation error is dependent on , the ratio of collision radius to taxi distance, with an order of 2.